# TopoFR: A Closer Look at Topology Alignment on Face Recognition

### v1

@@ Abstract
The field of face recognition (FR) has undergone significant advancements with the rise of deep learning. Recently, the success of unsupervised learning and graph neural networks has demonstrated the effectiveness of data structure information. Considering that the FR task can leverage large-scale training data, which intrinsically contains significant structure information, we aim to investigate how to encode such critical structure information into the latent space. As revealed from our observations, directly aligning the structure information between the input and latent spaces inevitably suffers from an overfitting problem, leading to a structure collapse phenomenon in the latent space. To address this problem, we propose TopoFR, a novel FR model that leverages a topological structure alignment strategy called PTSA and a hard sample mining strategy named SDE. Concretely, PTSA uses persistent homology to align the topological structures of the input and latent spaces, effectively preserving the structure information and improving the generalization performance of FR model. To mitigate the impact of hard samples on the latent space structure, SDE accurately identifies hard samples by automatically computing structure damage score (SDS) for each sample, and directs the model to prioritize optimizing these samples. Experimental results on popular face benchmarks demonstrate the superiority of our TopoFR over the state-of-the-art methods. Code and models are available at: https://github.com/modelscope/facechain/tree/main/face_module/TopoFR.
* Equal Contribution, $\dagger$ Corresponding Author.
@@ 1 Introduction
Face recognition (FR) is a critical biometric authentication technique that is widely applied in various applications. In recent years, convolutional neural networks (CNNs) have achieved remarkable success in FR task, thanks to their powerful ability to autonomously extract face features from images. Existing studies on FR primarily focuses on constructing more discriminative face features by developing margin-based loss functions [1, 2, 3, 4, 5] and powerful network architectures [6, 7, 8, 9]. Recently, the success of unsupervised learning [10, 11, 12, 13, 14] and graph neural networks [15, 16, 17] has demonstrated the importance of data structure information in improving model generalization. However, to the best of our knowledge, how to effectively mine the potential structure information in large-scale face data has not investigated. Thus, in this paper, we extend our interests on building a cutting-edge FR framework through exploiting such powerful and substantial structure information.
@@
Figure 1: We sample 1000 (a), 5000 (b), 10000 (c) and 100000 (d) face images from the MS1MV2 dataset respectively, and compute their persistence diagrams using PH, where $H_{j}$ represents the $j$-th dimension homology. Persistence diagram [18] is a mathematical tool to describe the topological structure of space, where the $j$-th dimension homology $H_{j}$ in persistence diagram represents the $j$-th dimension hole in space. In topology theory, if the number of high-dimensional holes in the space is more, then the underlying topological structure of the space is more complex [19]. As shown in Figure 1(a)-1(d), as the amount of face data increases, the persistence diagram of the input space contains more and more high-dimensional holes (e.g., $H_{3}$ and $H_{4}$). Therefore, this phenomenon demonstrates a growing complexity in the topological structure of the input space.
@@
First, we use Persistent Homology (PH) [20, 21], a mathematical tool used in topological data analysis [22] to capture the underlying topological structure of complex point clouds, to investigate the evolution trend of structure information in existing FR framework and illustrate 3 interesting observations: (1) as the amount of data increases, the topological structure of the input space becomes more and more complex, as verified in Figures 1a-1d; (2) as the amount of data increases, the topological structure discrepancy between the input space and the latent space becomes increasingly larger, as verified in Figure 2a; (3) The results in Figure 2b demonstrate that as the depth of the network increases, the topological structure discrepancy becomes progressively smaller. This finding also provides an explanation for why models with more complex structure achieve higher FR accuracy. Based on the above observations, we can infer that in FR tasks with large-scale datasets, the structure of face data will be severely destroyed during training, which limits the generalization ability of FR models in practical application scenarios. To this end, we propose to improve the generalization performance of FR models by preserving the structure information.
@@
However, we experimentally find that directly using PH to align the topological structures of the input space and the latent space may cause the model to suffer from structure collapse phenomenon. Concretely, under this experimental setting, we have 2 following quantitative results: (1) As shown in Figure 2c, the topological structure discrepancy drops to 0 dramatically during early training. (2) As depicted in Figure 2d, when evaluating on the IJB-C benchmark [23], there exists a significant structure information gap between the input space and the latent space. These typical overfitting phenomena indicate the latent space fails to preserve the structure information of input space accurately.
@@
Moreover, in practical FR scenarios, the training dataset typically includes some low-quality face samples (i.e., hard samples) that are prone to being encoded into abnormal positions close to the decision boundary in the latent space [24, 6, 25, 26], significantly destroying the topological structure of the latent space and affecting the alignment of structure. To address this issue, we propose a novel hard sample mining strategy named Structure Damage Estimation (SDE). SDE adaptively assigns structure damage score (SDS) to each sample based on its prediction uncertainty and prediction probability. By prioritizing the optimization of hard samples with significant structure damage, SDE can gradually guide these samples back to their reasonable positions, thereby improving the generalization ability of FR model.
Figure 2: (a): We investigate the relationship between the amount of data and the topological structure discrepancy by employing ResNet-50 ArcFace model [1] to perform inferences on MS1MV2 training set. Inferences are conducted for 1000 iterations with batch sizes of 256, 1024, and 2048, respectively. Histograms are used to approximate these discrepancy distributions. (b): We investigate the relationship between the network depth and the topological structure discrepancy by performing inference on MS1MV2 training set (batch size=128) using ArcFace models with different backbones. (c): We investigate the trend of topological structure discrepancy during training (batch size=128) and found that i) directly using PH to align the topological structures will cause the discrepancy to drops to 0 dramatically; ii) whereas using our PTSA strategy promotes a smooth convergence of structure discrepancy. (d): Aligning the topological structures directly using PH will lead to significant discrepancy when evaluating on IJB-C benchmark. Our PTSA strategy effectively mitigates this overfitting issue, resulting in smaller structure discrepancy during evaluation.
@@
3) Experimental results show that the proposed method outperforms SOTA methods on various face benchmarks. Notably, our TopoFR has secured the second place in the ICCV21 MFR-Ongoing challenge [27] until the submission of this work (May 22 ’24, academic track): http://iccv21-mfr.com/#/leaderboard/academic, indicating the robustness and generalization of our method.
@@ 2 Related Works
Face Recognition (FR). Convolutional Neural Networks (CNNs) have achieved remarkable advancements in tasks related to facial recognition  [28, 29, 30, 1, 31]. Notably, the extraction of robust deep facial embeddings has raised considerable interest within the research community. Among them, CNNs framework are representative methods, using two primary methods: metric learning-based and margin-based softmax approaches. The former utilizes loss functions like Triplet loss [7], Tuplet loss [32], and Center loss [33] to learn discriminative face features, while the latter aims to incorporate margin penalty into the softmax loss framework, including methods such as ArcFace [1], CosFace [2], AM-softmax [34], and SphereFace [35]. Recent studies have explored various techniques, including adaptive parameters [3, 5], mining [4, 36, 37], learning acceleration [38, 39, 40], vision transformer architecture [9, 8], and data uncertainty [25, 6, 26] to further enhance models’ performance on large-scale datasets.
Persistent Homology (PH). Over the past decade, PH has shown significant advantages in multiple various such as signal processing[41], video analysis [42], neuroscience [43, 44], disease diagnosis [45] and evaluation of embedding strategies [46, 47]. In the field of machine learning, some studies [48, 49, 50] have shown that integrating topological representations into network can enhance model’s recognition/segmentation performance. [51] proposes a topology distance for the evaluation of GANs.
@@ 3 Background: Persistent Homology
PH is a computational topology method that quantifies the changes in the topological invariants of a Vietoris-Rips complex as a scale parameter $\rho$ is varied. In this section, we introduce some key concepts of PH. Further details on PH can be found in Refs. [20, 52].
@@
Vietoris-Rips Complex. The Vietoris-Rips complex [53] is a special simplicial complex constructed from a set of points in a metric space, and it can be used to approximate the topology of the underlying space. For $0\leq\rho<\infty$, we represent the Vietoris-Rips complex of point cloud $\mathcal{X}$ at scale $\rho$ as $\mathcal{V}_{\rho}(\mathcal{X})$, which contains all simplices (i.e., subsets) of $\mathcal{X}$, and each component of point cloud $\mathcal{X}$ satisfies a distance constraint: $\mu(x_{i},x_{j})\leq\rho$ for any $i,j$. Moreover, the Vietoris-Rips complex satisfies a nesting relation, i.e., $\mathcal{V}_{\rho_{i}}\subseteq\mathcal{V}_{\rho_{j}}$ for any $\rho_{i}\leq\rho_{j}$, which allows us to track the evolution progress of simplical complex as the scale $\rho$ increases. It is worth noting that $\mathcal{V}_{\rho}(\mathcal{X})$ and $\mathcal{V}_{\rho}(\mathcal{M})$ are equivalent because constructing the Vietoris-Rips complex only requires distance.
Homology Group. The homology group [54] is an algebraic structures that analyzes the topological features of a simplicial complex in different dimension $j$, such as connected components ($H_{0}$), cycles ($H_{1}$), voids ($H_{2}$), and higher-dimensional features ($H_{j},j\geq 3$). By tracking the changes in topological features $(H_{j})$ of the Vietoris-Rips complex as the scale $\rho$ increases, it is possible to gain insight into the multi-scale topological information of the underlying space.
Persistence Diagram and Persistence Pairing [55]. The persistence diagram $\mathcal{D}$ is a multi-set of points $(b,d)$ in the Cartesian plane $\mathbb{R}^{2}$, which encodes information about the lifespan of topological features. Specially, it summarizes the birth time $b$ and death time $d$ of each topological feature, where birth time $b$ signifies the scale at which the feature is created and death time $d$ refers to the scale at which it is destroyed. The persistence pairing $\gamma$ contains indices $(i,j)$ corresponding to simplices $r_{i},r_{j}\in\mathcal{V}_{\rho}(\mathcal{X})$ that create and destroy the topological features identified by $(b,d)\in\mathcal{D}$, respectively.
@@ 4.1 Perturbation-guided Topological Structure Alignment
Random Structure Perturbation (RSP). PTSA first utilizes the RSP mechanism to randomly perturb the structure of the latent space. Specially, it utilizes a data augmentation list $\mathcal{A}=\left\{\mathcal{A}_{1},\mathcal{A}_{2},\mathcal{A}_{3},\mathcal{A}_{4}\right\}$ that includes four common data augmentation operations, namely Random Erasing $\mathcal{A}_{1}$, GaussianBlur $\mathcal{A}_{2}$, Grayscale $\mathcal{A}_{3}$ and ColorJitter $\mathcal{A}_{4}$. For each training sample $x_{i}$, RSP will randomly select an operation $\mathcal{A}_{r}$ from $\mathcal{A}$ to perturb it, i.e., $\widetilde{x}_{i}=\mathcal{A}_{r}(x_{i})$. Then the perturbed sample $\widetilde{x}_{i}$ will be fed into the model for supervised learning, which effectively increases the structure diversity of the latent space. In our model, we adopt ArcFace loss [1] as the basic classification loss:
@@
Invariant Structure Alignment (ISA). Given a mini-batch of original training samples $\mathcal{X}=\left\{x_{i}\right\}_{i=1}^{n}$, we denote the perturbed batch samples as $\widetilde{\mathcal{X}}=\left\{\widetilde{x}_{i}\right\}_{i=1}^{n}$. For the perturbed samples, we denote the latent features extracted by $\mathcal{F}$ as $\widetilde{\mathcal{Z}}=\left\{\widetilde{f}_{i}\right\}_{i=1}^{n}$. During forward propagation, we can construct the Vietoris-Rips complexes $\mathcal{V}_{\rho}(\mathcal{X})$ and $\mathcal{V}_{\rho}(\widetilde{\mathcal{Z}})$ for point clouds $\mathcal{X}$ and $\widetilde{\mathcal{Z}}$ respectively, based on their respective pairwise distance matrix $\mathcal{M}^{\mathcal{X}}$and $\mathcal{M}^{\widetilde{\mathcal{Z}}}$. Then we can utilize persistent homology to analyze the topological structures of $\mathcal{V}_{\rho}(\mathcal{X})$ and $\mathcal{V}_{\rho}(\widetilde{\mathcal{Z}})$, and obtain their corresponding persistence diagrams $\left\{\mathcal{D}^{\mathcal{X}},\mathcal{D}^{\widetilde{\mathcal{Z}}}\right\}$ and persistence pairings $\left\{\gamma^{\mathcal{X}},\gamma^{\widetilde{\mathcal{Z}}}\right\}$, respectively.
Ideally, no matter how the face image is perturbed, the position of the encoded face feature in the latent space should remain unchanged, and the topological structure of the perturbed latent space should also be consistent with the original input space. To this end, we choose to align the original input space $\mathcal{X}$ with the perturbed latent space $\mathcal{\widetilde{Z}}$ to achieve this goal. Prior studies usually utilize bottleneck distance or Wasserstein distance to measure the topological structure discrepancy [56, 18] between two spaces by comparing the differences in persistence diagrams. However, these two metrics are sensitive to outliers in persistence diagrams [51, 57] and will significantly increase models’ training time, rendering them unsuitable for FR tasks with extremely large-scale datasets, as verified in Table 8 in the Appendix.
To mitigate this issue, we turn to retrieve the persistence diagrams values by subsetting the corresponding pairwise distance matrix with edge indices provided by the persistence pairings [58, 59, 60], i.e., $\mathcal{D}^{\mathcal{X}}\simeq\mathcal{M}^{\mathcal{X}}[\gamma^{\mathcal{X}}]$ and $\mathcal{D}^{\mathcal{Z}}\simeq\mathcal{M}^{\widetilde{\mathcal{Z}}}[\gamma^{\widetilde{\mathcal{Z}}}]$. By comparing the difference between two topologically relevant distance matrices from both spaces, we can quickly and stably compute the discrepancy between their persistence diagrams, providing an efficient solution for structure alignment of FR models driven by large-scale datasets. We formulate the ISA loss as follows:
@@
Notably, in the field of FR, most existing works do not include any data augmentation operations, as this would introduce more unidentifiable face images (i.e., destroying the fidelity of each face), which generally hurts the FR model’s generalization ability, as verified in Refs. [3, 61]. In this work, we do not employ these data augmentations to simply augment data scale. Instead, we use them to increase the latent space’s structure diversity, effectively addressing the structure collapse problem. As a result, our PTSA strategy can reap the benefits of data augmentations while mitigating their potential negative effects (see Figure 2, Table 3, and Figure 5 for further analysis.).

### v2

@@ Abstract
The field of face recognition (FR) has undergone significant advancements with the rise of deep learning. Recently, the success of unsupervised learning and graph neural networks has demonstrated the effectiveness of data structure information. Considering that the FR task can leverage large-scale training data, which intrinsically contains significant structure information, we aim to investigate how to encode such critical structure information into the latent space. As revealed from our observations, directly aligning the structure information between the input and latent spaces inevitably suffers from an overfitting problem, leading to a structure collapse phenomenon in the latent space. To address this problem, we propose TopoFR, a novel FR model that leverages a topological structure alignment strategy called PTSA and a hard sample mining strategy named SDE. Concretely, PTSA uses persistent homology to align the topological structures of the input and latent spaces, effectively preserving the structure information and improving the generalization performance of FR model. To mitigate the impact of hard samples on the latent space structure, SDE accurately identifies hard samples by automatically computing structure damage score (SDS) for each sample, and directs the model to prioritize optimizing these samples. Experimental results on popular face benchmarks demonstrate the superiority of our TopoFR over the state-of-the-art methods. Code and models are available at: https://github.com/modelscope/facechain/tree/main/face_module/TopoFR. * Equal Contribution, $\dagger$ Corresponding Author.
@@ 1 Introduction
Face recognition (FR) is a critical biometric authentication technique that is widely applied in various applications. In recent years, convolutional neural networks (CNNs) have achieved remarkable success in FR task, thanks to their powerful ability to autonomously extract face features from images. Existing studies on FR primarily focuses on constructing more discriminative face features by developing margin-based loss functions deng2019arcface ; wang2018cosface ; kim2022adaface ; huang2020curricularface ; meng2021magface and powerful network architectures chang2020data ; schroff2015facenet ; zhong2021face ; dan2023transface . Recently, the success of unsupervised learning zhu2022manifold ; wu2023manifold ; jiang2021decentralized ; lopes2023self ; dan2023homda and graph neural networks dan2024hogda ; wong2021persistent ; wu2023difformer has demonstrated the importance of data structure information in improving model generalization. However, to the best of our knowledge, how to effectively mine the potential structure information in large-scale face data has not investigated. Thus, in this paper, we extend our interests on building a cutting-edge FR framework through exploiting such powerful and substantial structure information.
@@
Figure 1: We sample 1000 (a), 5000 (b), 10000 (c) and 100000 (d) face images from the MS1MV2 dataset respectively, and compute their persistence diagrams using PH, where $H_{j}$ represents the $j$-th dimension homology. Persistence diagram mileyko2011probability is a mathematical tool to describe the topological structure of space, where the $j$-th dimension homology $H_{j}$ in persistence diagram represents the $j$-th dimension hole in space. In topology theory, if the number of high-dimensional holes in the space is more, then the underlying topological structure of the space is more complex zomorodian2004computing . As shown in Figure 1(a)-1(d), as the amount of face data increases, the persistence diagram of the input space contains more and more high-dimensional holes (e.g., $H_{3}$ and $H_{4}$). Therefore, this phenomenon demonstrates a growing complexity in the topological structure of the input space.
@@
First, we use Persistent Homology (PH) edelsbrunner2008persistent ; barannikov1994framed , a mathematical tool used in topological data analysis carlsson2009topology to capture the underlying topological structure of complex point clouds, to investigate the evolution trend of structure information in existing FR framework and illustrate 3 interesting observations: (1) as the amount of data increases, the topological structure of the input space becomes more and more complex, as verified in Figures 1a-1d; (2) as the amount of data increases, the topological structure discrepancy between the input space and the latent space becomes increasingly larger, as verified in Figure 2a; (3) The results in Figure 2b demonstrate that as the depth of the network increases, the topological structure discrepancy becomes progressively smaller. This finding also provides an explanation for why models with more complex structure achieve higher FR accuracy. Based on the above observations, we can infer that in FR tasks with large-scale datasets, the structure of face data will be severely destroyed during training, which limits the generalization ability of FR models in practical application scenarios. To this end, we propose to improve the generalization performance of FR models by preserving the structure information.
@@
However, we experimentally find that directly using PH to align the topological structures of the input space and the latent space may cause the model to suffer from structure collapse phenomenon. Concretely, under this experimental setting, we have 2 following quantitative results: (1) As shown in Figure 2c, the topological structure discrepancy drops to 0 dramatically during early training. (2) As depicted in Figure 2d, when evaluating on the IJB-C benchmark maze2018iarpa , there exists a significant structure information gap between the input space and the latent space. These typical overfitting phenomena indicate the latent space fails to preserve the structure information of input space accurately.
@@
Moreover, in practical FR scenarios, the training dataset typically includes some low-quality face samples (i.e., hard samples) that are prone to being encoded into abnormal positions close to the decision boundary in the latent space dan2023uncertainty ; chang2020data ; li2021spherical ; shi2019probabilistic , significantly destroying the topological structure of the latent space and affecting the alignment of structure. To address this issue, we propose a novel hard sample mining strategy named Structure Damage Estimation (SDE). SDE adaptively assigns structure damage score (SDS) to each sample based on its prediction uncertainty and prediction probability. By prioritizing the optimization of hard samples with significant structure damage, SDE can gradually guide these samples back to their reasonable positions, thereby improving the generalization ability of FR model.
Figure 2: (a): We investigate the relationship between the amount of data and the topological structure discrepancy by employing ResNet-50 ArcFace model deng2019arcface to perform inferences on MS1MV2 training set. Inferences are conducted for 1000 iterations with batch sizes of 256, 1024, and 2048, respectively. Histograms are used to approximate these discrepancy distributions. (b): We investigate the relationship between the network depth and the topological structure discrepancy by performing inference on MS1MV2 training set (batch size=128) using ArcFace models with different backbones. (c): We investigate the trend of topological structure discrepancy during training (batch size=128) and found that i) directly using PH to align the topological structures will cause the discrepancy to drops to 0 dramatically; ii) whereas using our PTSA strategy promotes a smooth convergence of structure discrepancy. (d): Aligning the topological structures directly using PH will lead to significant discrepancy when evaluating on IJB-C benchmark. Our PTSA strategy effectively mitigates this overfitting issue, resulting in smaller structure discrepancy during evaluation.
@@
3) Experimental results show that the proposed method outperforms SOTA methods on various face benchmarks. Notably, our TopoFR has secured the second place in the ICCV21 MFR-Ongoing challenge deng2021masked until the submission of this work (May 22 ’24, academic track): http://iccv21-mfr.com/#/leaderboard/academic, indicating the robustness and generalization of our method.
@@ 2 Related Works
Face Recognition (FR). Convolutional Neural Networks (CNNs) liu2024tolerant ; liu2024envisioning have achieved remarkable advancements in tasks related to facial recognition  liu2022mogface ; liu2022damofd ; liu2020hambox ; deng2019arcface ; liu2020bfbox ; deng2020retinaface . Notably, the extraction of robust deep facial embeddings has raised considerable interest within the research community. Among them, CNNs framework are representative methods, using two primary methods: metric learning-based and margin-based softmax approaches. The former utilizes loss functions like Triplet loss schroff2015facenet , Tuplet loss sohn2016improved , and Center loss wen2016discriminative to learn discriminative face features, while the latter aims to incorporate margin penalty into the softmax loss framework, including methods such as ArcFace deng2019arcface , CosFace wang2018cosface , AM-softmax wang2018additive , and SphereFace liu2017sphereface . Recent studies have explored various techniques, including adaptive parameters kim2022adaface ; meng2021magface , mining huang2020curricularface ; xu2021consistent ; deng2020sub , learning acceleration an2021partial ; li2021virtual ; an2022killing , vision transformer architecture dan2023transface ; zhong2021face , and data uncertainty li2021spherical ; chang2020data ; shi2019probabilistic to further enhance models’ performance on large-scale datasets.
Persistent Homology (PH). Over the past decade, PH has shown significant advantages in multiple various such as signal processingperea2015sliding , video analysis tralie2018quasi ; liu2024omniclip , neuroscience singh2008topological ; dabaghian2012topological , disease diagnosis chung2018topological and evaluation of embedding strategies rieck2015persistent ; rieck2017agreement . In the field of machine learning, some studies clough2020topological ; khramtsova2022rethinking ; venkataraman2016persistent have shown that integrating topological representations into network can enhance model’s recognition/segmentation performance. horak2021topology proposes a topology distance for the evaluation of GANs.
@@ 3 Background: Persistent Homology
PH is a computational topology method that quantifies the changes in the topological invariants of a Vietoris-Rips complex as a scale parameter $\rho$ is varied. In this section, we introduce some key concepts of PH. Further details on PH can be found in Refs. edelsbrunner2008persistent ; edelsbrunner2000topological .
@@
Vietoris-Rips Complex. The Vietoris-Rips complex vietoris1927hoheren is a special simplicial complex constructed from a set of points in a metric space, and it can be used to approximate the topology of the underlying space. For $0\leq\rho<\infty$, we represent the Vietoris-Rips complex of point cloud $\mathcal{X}$ at scale $\rho$ as $\mathcal{V}_{\rho}(\mathcal{X})$, which contains all simplices (i.e., subsets) of $\mathcal{X}$, and each component of point cloud $\mathcal{X}$ satisfies a distance constraint: $\mu(x_{i},x_{j})\leq\rho$ for any $i,j$. Moreover, the Vietoris-Rips complex satisfies a nesting relation, i.e., $\mathcal{V}_{\rho_{i}}\subseteq\mathcal{V}_{\rho_{j}}$ for any $\rho_{i}\leq\rho_{j}$, which allows us to track the evolution progress of simplical complex as the scale $\rho$ increases. It is worth noting that $\mathcal{V}_{\rho}(\mathcal{X})$ and $\mathcal{V}_{\rho}(\mathcal{M})$ are equivalent because constructing the Vietoris-Rips complex only requires distance.
Homology Group. The homology group edelsbrunner2013persistent is an algebraic structures that analyzes the topological features of a simplicial complex in different dimension $j$, such as connected components ($H_{0}$), cycles ($H_{1}$), voids ($H_{2}$), and higher-dimensional features ($H_{j},j\geq 3$). By tracking the changes in topological features $(H_{j})$ of the Vietoris-Rips complex as the scale $\rho$ increases, it is possible to gain insight into the multi-scale topological information of the underlying space.
Persistence Diagram and Persistence Pairing cohen2005stability . The persistence diagram $\mathcal{D}$ is a multi-set of points $(b,d)$ in the Cartesian plane $\mathbb{R}^{2}$, which encodes information about the lifespan of topological features. Specially, it summarizes the birth time $b$ and death time $d$ of each topological feature, where birth time $b$ signifies the scale at which the feature is created and death time $d$ refers to the scale at which it is destroyed. The persistence pairing $\gamma$ contains indices $(i,j)$ corresponding to simplices $r_{i},r_{j}\in\mathcal{V}_{\rho}(\mathcal{X})$ that create and destroy the topological features identified by $(b,d)\in\mathcal{D}$, respectively.
@@ 4.1 Perturbation-guided Topological Structure Alignment
Random Structure Perturbation (RSP). PTSA first utilizes the RSP mechanism to randomly perturb the structure of the latent space. Specially, it utilizes a data augmentation list $\mathcal{A}=\left\{\mathcal{A}_{1},\mathcal{A}_{2},\mathcal{A}_{3},\mathcal{A}_{4}\right\}$ that includes four common data augmentation operations, namely Random Erasing $\mathcal{A}_{1}$, GaussianBlur $\mathcal{A}_{2}$, Grayscale $\mathcal{A}_{3}$ and ColorJitter $\mathcal{A}_{4}$. For each training sample $x_{i}$, RSP will randomly select an operation $\mathcal{A}_{r}$ from $\mathcal{A}$ to perturb it, i.e., $\widetilde{x}_{i}=\mathcal{A}_{r}(x_{i})$. Then the perturbed sample $\widetilde{x}_{i}$ will be fed into the model for supervised learning, which effectively increases the structure diversity of the latent space. In our model, we adopt ArcFace loss deng2019arcface as the basic classification loss:
@@
[…]
Ideally, no matter how the face image is perturbed, the position of the encoded face feature in the latent space should remain unchanged, and the topological structure of the perturbed latent space should also be consistent with the original input space. To this end, we choose to align the original input space $\mathcal{X}$ with the perturbed latent space $\mathcal{\widetilde{Z}}$ to achieve this goal. Prior studies usually utilize bottleneck distance or Wasserstein distance to measure the topological structure discrepancy turner2014frechet ; mileyko2011probability between two spaces by comparing the differences in persistence diagrams. However, these two metrics are sensitive to outliers in persistence diagrams horak2021topology ; bertsekas1981new and will significantly increase models’ training time, rendering them unsuitable for FR tasks with extremely large-scale datasets, as verified in Table 8 in the Appendix.
To mitigate this issue, we turn to retrieve the persistence diagrams values by subsetting the corresponding pairwise distance matrix with edge indices provided by the persistence pairings zomorodian2010fast ; cohen2009extending ; moor2020topological , i.e., $\mathcal{D}^{\mathcal{X}}\simeq\mathcal{M}^{\mathcal{X}}[\gamma^{\mathcal{X}}]$ and $\mathcal{D}^{\mathcal{Z}}\simeq\mathcal{M}^{\widetilde{\mathcal{Z}}}[\gamma^{\widetilde{\mathcal{Z}}}$]. By comparing the difference between two topologically relevant distance matrices from both spaces, we can quickly and stably compute the discrepancy between their persistence diagrams, providing an efficient solution for structure alignment of FR models driven by large-scale datasets. We formulate the ISA loss as follows:
@@
Notably, in the field of FR, most existing works do not include any data augmentation operations, as this would introduce more unidentifiable face images (i.e., destroying the fidelity of each face), which generally hurts the FR model’s generalization ability, as verified in Refs. kim2022adaface ; shi2020towards . In this work, we do not employ these data augmentations to simply augment data scale. Instead, we use them to increase the latent space’s structure diversity, effectively addressing the structure collapse problem. As a result, our PTSA strategy can reap the benefits of data augmentations while mitigating their potential negative effects (see Figure 2, Table 3, and Figure 5 for further analysis.).